OpenStudy (anonymous):

The product of 4 and a number is 18 more than twice the number. What is the number?

OpenStudy (anonymous):

@Nnesah

OpenStudy (ajspeller):

The product means to multiply we dont know the number so call it x the word is usually means = 18 more means addition twice the number means 2 times x 4x = 18 + 2x The rest is up to you.

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (daniellelovee):

4x=18+2x since x is the number and 2x is twice the number

OpenStudy (daniellelovee):

now work backward PEMDAS with same variables therefore 4x-2x=2x so 2x=18 and you probably know what to do from there :)

OpenStudy (anonymous):

I don't know really I'm stuck

OpenStudy (daniellelovee):

ok hold on

OpenStudy (daniellelovee):



Nnesha (nnesha):

:D

OpenStudy (anonymous):

lol

OpenStudy (daniellelovee):

so makes sense?

OpenStudy (anonymous):

Yep thanks ^_^

OpenStudy (daniellelovee):

np :)